# Determinantal random subgraphs

### v3

@@ 11.2.1. Vertices as parameters
The quantity $\det(q_{i}(V_{j}))_{1\leqslant i,j\leqslant k}$ is nonzero, equal to ±$\pm 1$, if and only if there is a bijection $\sigma$ of $\{1,\ldots,k\}$ such that for all $i\in\{1,\ldots,k\}$, we have $v_{i}\in V_{\sigma(i)}$.
@@
Conversely, suppose we have such a bijection $\sigma$. Observe that we then have $q_{i}(V_{\sigma(i)})=1$ and $q_{i}(V_{j})=0$ if $j\neq\sigma(i)$, which shows that the matrix of which we are taking the determinant is the permutation matrix associated with $\sigma$. In particular, this determinant is nonzero and equal to ±$\pm 1$. This concludes the proof. ∎
@@ 11.2.2. Rooted spanning forests
This is a special case of Lemma 11.1 above. However, we give a more pedestrian and ‘graphical’ proof. Let $f\in\Omega^{0}(\mathsf{G})$. Note that for any path $c$ connecting vertex $u$ to vertex $v$, the result of a telescoping sum gives $df(c)=f(v)-f(u)$. Hence, $df(c_{i})=0$ for all $1\leqslant i\leqslant k$, if and only if $f(v_{i})=f(v_{0})$ for all $1\leqslant i\leqslant k$. This concludes the proof. ∎
@@ 11.2.4. Identity between two Green functions
(93) $\mathsf{A}_{k}({\sf c},\underline{x})=\det\Delta_{\widehat{\{v_{0},v_{1},\ldots,v_{k}\}}}^{\widehat{\{v_{0},v_{1},\ldots,v_{k}\}}}\,.$
@@ 11.3.1. Various Laplacian operators and their spectra
In the language of [Ken11, Kas15, KL21, KL20, KL22, KL25], this is the covariant Laplacian associated to a connection on the trivial vector bundle of rank $1$ with connection equal to $-1$ on all edges. It is a self-adjoint nonnegative operator on $\Omega^{0}(\mathsf{G})$.
@@
Hence $2\mu$ is a nonzero eigenvalue in $\mathcal{L}$ since $f\neq 0$ (because $\phi\notin\ker\mathsf{k}=\ker d^{*}$). Since $\chi^{-1/2}d^{*}$ is $1$-to-$1$ on $\mathop{\mathrm{im}}d$, by taking $\phi$ spanning a basis of $\mathop{\mathrm{im}}d$ consisting in eigenvectors of $\mathsf{k}$, we find $n$ corresponding eigenvectors for $\mathcal{L}$. Since $n=\mathrm{rank}(\mathcal{L})$, we have found all nonzero eigenvalues and the claim is proved. ∎
@@ 11.3.2. The Laplacian spanning forest
See Figure 6 (left) for a sample of this distribution. Since the connected components of this forest seem rather small, it is tempting to look empirically at the process whose kernel is $\mathsf{k}^{a}$ where the exponent $a$ is chosen between 00 and $1$. See Figure 6 (right).
@@ 11.3.4. Gap probability and bipartiteness condition
We note that $\mathsf{k}$ has eigenvalue $1$ if and only if the operator $2\chi-d^{*}d$ has eigenvalue 00, but since this operator is equal to the twisted Laplacian $\Delta_{(-1)}$ with connection $-1$ on all edges, we know from [For93] (see also [KL20]) that it has a nontrivial kernel if and only if for all simple cycles, the product of weights is equal to $1$ (the signed graph is said to be balanced). Since for any cycle, this product is $(-1)^{\text{length of cycle}}$, $\mathsf{k}$ has eigenvalue $1$ if and only if all cycles are of even lengths, which means that $\mathsf{G}$ is bipartite. ∎
@@ 12.1.1. Random $g$-quasitrees
Our observation is that [LLY25, Theorem 1.2] can be interpreted as an equality of bilinear forms on $^{k}\Omega^{1}(\mathsf{G})$: one defined via the period matrix, and the other via a ratio of graph polynomials. Interestingly, the identities of our Proposition 8.2, taken for all $k\in\{1,\ldots,2g\}$, yield the corresponding equalities of quadratic forms, which by polarization encode the same information.
@@
This crossing of perspectives yields new insight. In particular, Section 7.1 of [LLY25] suggests a geometric refinement of harmonic 1-forms that we had not mentioned in the first version of our paper. Indeed it points to the importance of Lagrangian subspaces (which are of dimension $g$) within the space of harmonic 1-forms (which has dimension $2g$). These are the subspaces of harmonic $1$-forms obtained from the span of $g$ $1$-forms dual to a set of disjoint, homologically independent cycles. This highlights a distinguished family of determinant probability measures on random $g$-quasitrees associated with a surface discretization. These are parametrized by the compact manifold (of dimension $g(g+1)/2$) of all Lagrangian subspaces, which is called the Lagrangian Grassmannian. See Figure 7 for an example.
Figure 7. The $2$-core of a determinantal $2$-quasitree associated with $1$-forms dual to the two meridians.
@@ 12.1.3. Three-dimensional cylinder case
Figure 9. Exact sample of the probability distribution from Theorem 12.1 in a square grid approximation of a thick walled pipe $M=[0,10]^{3}\setminus[3,4]\times[0,10]\times[2,3]$. The cycle winds twice around the central tube. The $1$-form generating $H_{1}$ chosen for the simulation is supported on the half-wall of green edges. Red edges show places where the cycle takes one of those green edges.
@@ 12.1.4. Laplacian spanning connected subgraph
[…]
@@ 13.1.2. Representing map and its kernel
=$\mathcal{I}=\big{\{}I\subseteq S:\text{the family }\{\partial(e_{i}):i\in I\}\text{ is linearly independent in }F\big{\}}.$
@@ 13.1.4. Union of a matroid with a uniform matroid
A subset $K$ of $S$ which contains a basis of $\mathcal{M}$ must have a cardinality of the form $n+k$ for some $k\in\{0,\ldots,b\}$, because $b=s-n$. For each such integer $k$, the set
@@ 13.3. The circuit-basis identities for linear matroids
Then in $^{k}E_{K}$,
@@
Moreover, in ($^{k}E_{K}\big{)}^{\otimes 2}$, we have
@@ 2. For every $T\in\mathcal{B}(\mathcal{M})$, we have
Now, $^{n}\mathop{\mathrm{im}}d$ is a line, generated by the unit norm vector $\eta$, so that $({\sf\Pi}^{\mathop{\mathrm{im}}d})^{\wedge n}e^{\star}_{T}=\big{\langle}\eta,e^{\star}_{T}\big{\rangle}\eta$. Therefore,
@@ 13.7. Conditional probability measures
${\mathcal{M}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$prob.
@@ 13.8. A geometric multilinear identity
In $^{k}E\otimes_{\mathbb{R}}\mathchoice{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{0.5pt}{$\scriptstyle\bigwedge$}}{\raisebox{0.2pt}{$\scriptscriptstyle\bigwedge$}}^{k}E$, the following equality holds:
@@
where the sum is over all pairs formed by a subset $K\in\mathcal{B}_{k}$ and a basis $T$ of $\mathcal{M}_{|K}$. Let us re-index this sum as a sum over pairs formed by a basis $T$ of $\mathcal{M}$ and a $k$-subset $I$ of $S$ disjoint from $T$, this set $I$ playing the role of $K\setminus T$. We find
@@
and since $({\mathsf{P}}_{T})^{\wedge k}(e^{\star}_{I})=0$ whenever the subset $I$ is not disjoint from the basis $T$, we can read the last sum as a double sum over all bases $T$ of $\mathcal{M}$ and all $k$-subsets $I$ of $S$. An application of the mean projection theorem (Proposition 13.7) concludes the proof. ∎
@@ 14.1.4. Twisted Green function minors
[…]
@@
We first note that each side of the equation we want to prove is a quadratic form in $q_{1}\wedge\ldots\wedge q_{k}$. Hence it is enough to prove it for an orthonormal basis of $^{k}\Omega^{0}(\mathsf{G})$, and then conclude by bilinearity. We will thus prove the formula for each collection of disjoint $k$ vertices $v_{i}$, letting $q_{i}=\mathbbm{1}_{v_{i}}$.
@@
is nonzero if and only if there exists a bijection $\sigma$ of $\{1,\ldots,k\}$ such that $v_{i}$ belongs to $T_{\sigma(i)}$, in which case, the determinant is the signature of this permutation matrix, equal to ±$\pm 1$. Hence in that case, $\mathsf{A}^{\rho}_{k}({\sf q},\underline{x})=\det(\Delta_{h})_{\widehat{W}}^{\widehat{W}}$. The result now follows by invoking (147). ∎
@@
(150) $^{k}{\sf\Pi}^{\mathop{\mathrm{im}}d_{h}}(\varphi)\rVert^{2}=\frac{\mathsf{A}^{\rho}_{k}({\sf q},\underline{x})}{\mathsf{B}_{\rho}(\underline{x})}\,.$
@@ 14.2.2. Perturbation of cycle-rooted spanning forests
As proved by Kenyon in [Ken11, Theorem 3], when we consider $\varepsilon\theta$ in place of $\theta$, then, letting $\varepsilon$ tend to 00, we have a family of determinantal measures on $\mathcal{U}(\mathsf{G})$ converging to a determinantal measure on $\mathcal{C}_{1}(\mathsf{G})$, which is precisely the measure described in Theorem 8.1 for the line $\Theta=\mathbb{K}\theta$. This convergence result is generalised to higher rank in [KL25]; see Section 14.3 for a preview.
@@ 14.3. Quantum spanning forests
Figure 13. The marginals of a rank-$2$ quantum spanning forest on a $15\times 15$ grid.
@@
In the case where the connection $h$ tends to a trivial one, the authors obtain a determinantal probability measure on $N$-tuples of connected subgraphs whose total first Betti number is $N$. The partition function of this probability measure is quite analogous to the one in the Symanzik case considered above in the case $k=N$. We wonder whether one can obtain the above measure as the limit of a richer probability distribution, or even as a certain marginal or conditioning of the quantum spanning forest.
@@ 15.1. Counting complexity, determinantal measures, and uniform sampling
Finding the number of elements of $\mathcal{U}(\mathsf{G})$ is known to be #$\#P$-hard [GN06, Section 3] [^13] The authors of that paper use a short two-step reduction to the counting problem of perfect matchings of a graph, which is known to be #$\#\mathsf{P}$-hard by a celebrated work of Valiant [Val79], where this computational complexity class was in fact introduced., and thus there can be no polynomial time computable formula for it. [^14] Bounds on the cardinality of $\mathcal{U}(\mathsf{G})$ in terms of that of $\mathcal{T}(\mathsf{G})$ were obtained in [GdMN05]. In particular, the uniform measure on $\mathcal{U}(\mathsf{G})$ cannot be determinantal, by [Lyo03, Corollary 5.5]. However, the uniform measure on $\mathcal{U}(\mathsf{G})$ may be sampled exactly in polynomial time, see [KK17, Theorem 1], [Kas15, Section 2.4], and [GJ21]. [^15] Similar variations on Wilson’s algorithm [Wil96] were proposed in [BBGJ07, GP14]. A general theory of partial rejection sampling was developped recently in [Jer21] which encompasses these as special cases. This yields a fully-polynomial approximation scheme (FPRAS) for enumerating $\mathcal{U}(\mathsf{G})$, as shown in [GJ21].
@@
Enumerating elements in $\mathcal{F}_{k}(\mathsf{G})$ can be done in polynomial time, using a combination of determinants (see [LC81], further simplified by [Myr92] and [KW16]). Similarly, according to [Cho92], the set $\mathcal{C}_{k}(\mathsf{G})$ can be enumerated in polynomial time. However, the enumerations of $\cup_{k\geqslant 0}\mathcal{F}_{k}(\mathsf{G})$ and $\cup_{k\geqslant 0}\mathcal{C}_{k}(\mathsf{G})$ are known to be impossible in polynomial time, as they are #$\#P$-hard evaluations of the Tutte polynomial of $\mathsf{G}$, see [Wel93, PB83]. In particular, the uniform measure on these sets is not determinantal in general. It is nevertheless conjectured that they satisfy a form of negative dependence, see [GW04].

### v4

@@ 11.2.1. Vertices as parameters
The quantity $\det(q_{i}(V_{j}))_{1\leqslant i,j\leqslant k}$ is nonzero, equal to $\pm 1$, if and only if there is a bijection $\sigma$ of $\{1,\ldots,k\}$ such that for all $i\in\{1,\ldots,k\}$, we have $v_{i}\in V_{\sigma(i)}$.
@@
Conversely, suppose we have such a bijection $\sigma$. Observe that we then have $q_{i}(V_{\sigma(i)})=1$ and $q_{i}(V_{j})=0$ if $j\neq\sigma(i)$, which shows that the matrix of which we are taking the determinant is the permutation matrix associated with $\sigma$. In particular, this determinant is nonzero and equal to $\pm 1$. This concludes the proof. ∎
@@ 11.2.2. Rooted spanning forests
This is a special case of Lemma 11.1 above. However, we give a more pedestrian and ‘graphical’ proof. Let $f\in\Omega^{0}(\mathsf{G})$. Note that for any path $c$ connecting vertex $u$ to vertex $v$, the result of a telescoping sum gives df(c)=f(v)f(u)df(c)=f(v)-f(u). Hence, $df(c_{i})=0$ for all $1\leqslant i\leqslant k$, if and only if $f(v_{i})=f(v_{0})$ for all $1\leqslant i\leqslant k$. This concludes the proof. ∎
@@ 11.2.4. Identity between two Green functions
[…]
@@ 11.3.1. Various Laplacian operators and their spectra
In the language of [Ken11, Kas15, KL21, KL20, KL22, KL25], this is the covariant Laplacian associated to a connection on the trivial vector bundle of rank $1$ with connection equal to 1-1 on all edges. It is a self-adjoint nonnegative operator on $\Omega^{0}(\mathsf{G})$.
@@
Hence $2\mu$ is a nonzero eigenvalue in $\mathcal{L}$ since $f\neq 0$ (because $\phi\notin\ker\mathsf{k}=\ker d^{*}$). Since $\chi^{-1/2}d^{*}$ is 11-to-11 on $\mathop{\mathrm{im}}d$, by taking $\phi$ spanning a basis of $\mathop{\mathrm{im}}d$ consisting in eigenvectors of $\mathsf{k}$, we find $n$ corresponding eigenvectors for $\mathcal{L}$. Since $n=\mathrm{rank}(\mathcal{L})$, we have found all nonzero eigenvalues and the claim is proved. ∎
@@ 11.3.2. The Laplacian spanning forest
See Figure 6 (left) for a sample of this distribution. Since the connected components of this forest seem rather small, it is tempting to look empirically at the process whose kernel is $\mathsf{k}^{a}$ where the exponent $a$ is chosen between 0 and $1$. See Figure 6 (right).
@@ 11.3.4. Gap probability and bipartiteness condition
We note that $\mathsf{k}$ has eigenvalue $1$ if and only if the operator $2\chi-d^{*}d$ has eigenvalue 0, but since this operator is equal to the twisted Laplacian $\Delta_{(-1)}$ with connection 1-1 on all edges, we know from [For93] (see also [KL20]) that it has a nontrivial kernel if and only if for all simple cycles, the product of weights is equal to $1$ (the signed graph is said to be balanced). Since for any cycle, this product is $(-1)^{\text{length of cycle}}$, $\mathsf{k}$ has eigenvalue $1$ if and only if all cycles are of even lengths, which means that $\mathsf{G}$ is bipartite. ∎
@@ 12.1.1. Random $g$-quasitrees
Our observation is that [LLY25, Theorem 1.2] can be interpreted as an equality of bilinear forms on $\mathchoice{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{0.5pt}{$\scriptstyle\bigwedge$}}{\raisebox{0.2pt}{$\scriptscriptstyle\bigwedge$}}^{k}\Omega^{1}(\mathsf{G})$: one defined via the period matrix, and the other via a ratio of graph polynomials. Interestingly, the identities of our Proposition 8.2, taken for all $k\in\{1,\ldots,2g\}$, yield the corresponding equalities of quadratic forms, which by polarization encode the same information.
@@
This crossing of perspectives yields new insight. In particular, Section 7.1 of [LLY25] suggests a geometric refinement of harmonic 1-forms that we had not mentioned in the first version of our paper. Indeed it points to the importance of Lagrangian subspaces (which are of dimension $g$) within the space of harmonic 1-forms (which has dimension $2g$). These are the subspaces of harmonic $1$-forms obtained from the span of $g$ 11-forms dual to a set of disjoint, homologically independent cycles. This highlights a distinguished family of determinant probability measures on random $g$-quasitrees associated with a surface discretization. These are parametrized by the compact manifold (of dimension $g(g+1)/2$) of all Lagrangian subspaces, which is called the Lagrangian Grassmannian. See Figure 7 for an example.
Figure 7. The 22-core of a determinantal 22-quasitree associated with 11-forms dual to the two meridians.
@@ 12.1.3. Three-dimensional cylinder case
Figure 9. Exact sample of the probability distribution from Theorem 12.1 in a square grid approximation of a thick walled pipe $M=[0,10]^{3}\setminus[3,4]\times[0,10]\times[2,3]$. The cycle winds twice around the central tube. The 11-form generating $H_{1}$ chosen for the simulation is supported on the half-wall of green edges. Red edges show places where the cycle takes one of those green edges.
@@ 12.1.4. Laplacian spanning connected subgraph
(115) $\Omega^{0}({\mathsf{X}})\,\overset{d}{\underset{d^{*}}{\begin{subarray}{c}\longrightarrow\\ \longleftarrow\end{subarray}}}\,\Omega^{1}(X)\,\overset{d}{\underset{d^{*}}{\begin{subarray}{c}\longrightarrow\\ \longleftarrow\end{subarray}}}\,\Omega^{2}(X)$
@@ 13.1.2. Representing map and its kernel
$\mathcal{I}=\big\{I\subseteq S:\text{the family }\{\partial(e_{i}):i\in I\}\text{ is linearly independent in }F\big\}.$
@@ 13.1.4. Union of a matroid with a uniform matroid
A subset $K$ of $S$ which contains a basis of $\mathcal{M}$ must have a cardinality of the form $n+k$ for some $k\in\{0,\ldots,b\}$, because b=snb=s-n. For each such integer $k$, the set
@@ 13.3. The circuit-basis identities for linear matroids
Then in $\mathchoice{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{0.5pt}{$\scriptstyle\bigwedge$}}{\raisebox{0.2pt}{$\scriptscriptstyle\bigwedge$}}^{k}E_{K}$,
@@
Moreover, in $\big(\mathchoice{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{0.5pt}{$\scriptstyle\bigwedge$}}{\raisebox{0.2pt}{$\scriptscriptstyle\bigwedge$}}^{k}E_{K}\big)^{\otimes 2}$, we have
@@ 2. For every $T\in\mathcal{B}(\mathcal{M})$, we have
Now, $\mathchoice{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{0.5pt}{$\scriptstyle\bigwedge$}}{\raisebox{0.2pt}{$\scriptscriptstyle\bigwedge$}}^{n}\mathop{\mathrm{im}}d$ is a line, generated by the unit norm vector $\eta$, so that $({\sf\Pi}^{\mathop{\mathrm{im}}d})^{\wedge n}e^{\star}_{T}=\big\langle\eta,e^{\star}_{T}\big\rangle\eta$. Therefore,
@@ 13.7. Conditional probability measures
[…]
@@ 13.8. A geometric multilinear identity
[…]
@@
where the sum is over all pairs formed by a subset $K\in\mathcal{B}_{k}$ and a basis $T$ of $\mathcal{M}_{|K}$. Let us re-index this sum as a sum over pairs formed by a basis $T$ of $\mathcal{M}$ and a $k$-subset II of $S$ disjoint from $T$, this set $I$ playing the role of $K\setminus T$. We find
@@
and since $({\mathsf{P}}_{T})^{\wedge k}(e^{\star}_{I})=0$ whenever the subset $I$ is not disjoint from the basis $T$, we can read the last sum as a double sum over all bases $T$ of $\mathcal{M}$ and all $k$-subsets II of $S$. An application of the mean projection theorem (Proposition 13.7) concludes the proof. ∎
@@ 14.1.4. Twisted Green function minors
Note that the right-hand side of (149) can be written in the language of the exterior algebra as $\langle q_{1}\wedge\ldots\wedge q_{k},\mathchoice{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{0.5pt}{$\scriptstyle\bigwedge$}}{\raisebox{0.2pt}{$\scriptscriptstyle\bigwedge$}}^{k}G_{h}(q_{1}\wedge\ldots\wedge q_{k})$ $\rangle$.
@@
We first note that each side of the equation we want to prove is a quadratic form in $q_{1}\wedge\ldots\wedge q_{k}$. Hence it is enough to prove it for an orthonormal basis of $\mathchoice{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{0.5pt}{$\scriptstyle\bigwedge$}}{\raisebox{0.2pt}{$\scriptscriptstyle\bigwedge$}}^{k}\Omega^{0}(\mathsf{G})$, and then conclude by bilinearity. We will thus prove the formula for each collection of disjoint $k$ vertices $v_{i}$, letting $q_{i}=\mathbbm{1}_{v_{i}}$.
@@
is nonzero if and only if there exists a bijection $\sigma$ of $\{1,\ldots,k\}$ such that $v_{i}$ belongs to $T_{\sigma(i)}$, in which case, the determinant is the signature of this permutation matrix, equal to $\pm 1$. Hence in that case, $\mathsf{A}^{\rho}_{k}({\sf q},\underline{x})=\det(\Delta_{h})_{\widehat{W}}^{\widehat{W}}$. The result now follows by invoking (147). ∎
@@
(150) $\lVert\mathchoice{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{0.5pt}{$\scriptstyle\bigwedge$}}{\raisebox{0.2pt}{$\scriptscriptstyle\bigwedge$}}^{k}{\sf\Pi}^{\mathop{\mathrm{im}}d_{h}}(\varphi)\rVert^{2}=\frac{\mathsf{A}^{\rho}_{k}({\sf q},\underline{x})}{\mathsf{B}_{\rho}(\underline{x})}\,.$
@@ 14.2.2. Perturbation of cycle-rooted spanning forests
As proved by Kenyon in [Ken11, Theorem 3], when we consider $\varepsilon\theta$ in place of $\theta$, then, letting $\varepsilon$ tend to 0, we have a family of determinantal measures on $\mathcal{U}(\mathsf{G})$ converging to a determinantal measure on $\mathcal{C}_{1}(\mathsf{G})$, which is precisely the measure described in Theorem 8.1 for the line $\Theta=\mathbb{K}\theta$. This convergence result is generalised to higher rank in [KL25]; see Section 14.3 for a preview.
@@ 14.3. Quantum spanning forests
Figure 13. The marginals of a rank-22 quantum spanning forest on a $15\times 15$ grid.
@@
In the case where the connection $h$ tends to a trivial one, we obtain a determinantal probability measure on $N$-tuples of connected subgraphs whose total first Betti number is $N$. The partition function of this probability measure is quite analogous to the one in the Symanzik case considered above in the case $k=N$. We wonder whether one can obtain the above measure as the limit of a richer probability distribution, or even as a certain marginal or conditioning of the quantum spanning forest.
@@ 15.1. Counting complexity, determinantal measures, and uniform sampling
Finding the number of elements of $\mathcal{U}(\mathsf{G})$ is known to be #P\#P-hard [GN06, Section 3] [^15] The authors of that paper use a short two-step reduction to the counting problem of perfect matchings of a graph, which is known to be $\mathsf{P}-$hard by a celebrated work of Valiant [Val79], where this computational complexity class was in fact introduced., and thus there can be no polynomial time computable formula for it. [^16] Bounds on the cardinality of $\mathcal{U}(\mathsf{G})$ in terms of that of $\mathcal{T}(\mathsf{G})$ were obtained in [GdMN05]. In particular, the uniform measure on $\mathcal{U}(\mathsf{G})$ cannot be determinantal, by [Lyo03, Corollary 5.5]. However, the uniform measure on $\mathcal{U}(\mathsf{G})$ may be sampled exactly in polynomial time, see [KK17, Theorem 1], [Kas15, Section 2.4], and [GJ21]. [^17] Similar variations on Wilson’s algorithm [Wil96] were proposed in [BBGJ07, GP14]. A general theory of partial rejection sampling was developped recently in [Jer21] which encompasses these as special cases. This yields a fully-polynomial approximation scheme (FPRAS) for enumerating $\mathcal{U}(\mathsf{G})$, as shown in [GJ21].
@@
Enumerating elements in $\mathcal{F}_{k}(\mathsf{G})$ can be done in polynomial time, using a combination of determinants (see [LC81], further simplified by [Myr92] and [KW16]). Similarly, according to [Cho92], the set $\mathcal{C}_{k}(\mathsf{G})$ can be enumerated in polynomial time. However, the enumerations of $\cup_{k\geqslant 0}\mathcal{F}_{k}(\mathsf{G})$ and $\cup_{k\geqslant 0}\mathcal{C}_{k}(\mathsf{G})$ are known to be impossible in polynomial time, as they are #P\#P-hard evaluations of the Tutte polynomial of $\mathsf{G}$, see [Wel93, PB83]. In particular, the uniform measure on these sets is not determinantal in general. It is nevertheless conjectured that they satisfy a form of negative dependence, see [GW04].